# G2D2: Gradient-guided Discrete Diffusion for image inverse problem solving

### v1

@@ 3.2 G2D2 based on star-shaped noise process
7:         Initialize: $\bm{\alpha}_{t}=\exp(\gamma\log\bm{\alpha}_{t+1}+(1-\gamma)\log\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t}))$
@@
11:     Sample $\mathbf{z}_{t-1}\sim p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})=\sum_{\mathbf{z}_{0}}q(\mathbf{z}_{t-1}|\mathbf{z}_{0})\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$
@@ 3.3 Implementation considerations
We use the Gumbel-Softmax trick (Jang et al., 2016; Maddison et al., 2016) to make the computation of the second term in (7) differentiable. At time step $t$, this process begins by generating Gumbel-Softmax samples using parameters of $\tilde{p}_{\bm{\alpha}}$ as follows: $\hat{z}_{0,i,k}=\mathrm{softmax}\left(\left(\log\alpha_{t,i,k}+g_{i,k}\right)/\tau\right)$, where $g_{i,k}$ represents samples drawn from the Gumbel distribution, and $\tau$ is the temperature parameter. This procedure generates a “soft” categorical sample for each dimension in $\mathbf{z}_{0}$, indicating the proportional selection of each codebook element. As these proportions correspond to the contribution rate of each codebook element, we construct $\mathbf{Z}_{\text{Gumbel}}\in\mathbb{R}^{d_{\mathbf{z}}\times d_{b}}$ as their weighted sum: $(\mathbf{Z}_{\text{Gumbel}})_{i}=\sum_{k=1}^{K}\hat{z}_{0,i,k}\mathbf{b}_{k}$. Finally, we pass $\mathbf{Z}_{\text{Gumbel}}$ through the decoder to obtain the image $\mathbf{x}_{0}=D(\mathbf{Z}_{\text{Gumbel}})$. By substituting this image into the likelihood function $q(\mathbf{y}|\mathbf{x}_{0})$, we have the differentiable objective with respect to the variational parameter $\bm{\alpha}_{t}$, enabling continuous optimization. For linear inverse problems, the objective function will include the term $\|\mathbf{y}-\mathbf{A}\mathbf{x}_{0}(\bm{\alpha}_{t})\|_{2}^{2}$, excluding the constant term derived from measurement noise.
@@
At time step $t$, we are required to optimize the variational parameter $\bm{\alpha}_{t}$. To expedite this process, we can leverage the optimized values from the previous time step as the initialization for the optimization process, effectively reducing the number of required optimization steps. To achieve this, we introduce a forgetting coefficient $\gamma$ and initialize $\bm{\alpha}_{t}$ through a weighted sum of the previous optimized variables and the prior model’s output in the logarithm domain, given by $\bm{\alpha}_{t}=\exp(\gamma\log\bm{\alpha}_{t+1}+(1-\gamma)\log\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t}))$.
@@ 3.4 Application of G2D2 to masked generative models
As discussed in (Zheng et al., 2024), mask-absorbing discrete diffusion models and masked generative models, such as MaskGIT (Chang et al., 2022), share a similar framework. Except for temporal conditioning, these models are nearly identical and are trained to approximate $q(\mathbf{z}_{0}|\mathbf{z}_{t})$. Therefore, G2D2 can be straightforwardly applied to masked generative models. We give an example of solving inverse problems using a masked generative model as a prior model for motion data in the following section.
@@ 4.1 Experimental setup
We evaluate G2D2 on inverse problems in image processing and compare it with other diffusion model-based inverse-problem-solving methods. We also demonstrate gradient-based guidance on a discrete-latent variable-based motion-domain generative model without additional training, showing the applicability of G2D2 to other domains.
@@
We conduct experiments on two tasks: (1) super-resolution (SR) and (2) Gaussian deblurring. For the SR task, the linear forward operator downscales the image by a factor of 4 using a bicubic resizer. For the Gaussian-deblurring task, we set the kernel size to $61\times 61$ with a Gaussian kernel standard deviation of 3.0. The measurements are obtained by applying the forward operator to the ground truth images normalized to the range $[-1,1]$, followed by the addition of Gaussian noise with a standard deviation of 0.05. As metrics, we use the learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018) score to measure perceptual proximity to the original image, and the peak signal-to-noise ratio (PSNR) to measure the closeness of the signal.
@@
Following previous studies, we use the ImageNet (Deng et al., 2009) and Flickr-Faces-HQ (FFHQ) (Karras et al., 2019) datasets. The size of both datasets is 256$\times$256. For comparison, we use a subset of 100 images from each validation set.
@@
We compare DPS (Chung et al., 2023), DDRM (Kawar et al., 2022), which use diffusion models trained in the pixel domain, and PSLD (Rout et al., 2023) and ReSample (Song et al., 2024), which use diffusion models trained in the latent space acquired from VAE (latent diffusion models) as baselines with G2D2.
@@
Regarding G2D2, for both the ImageNet and FFHQ experiments, we use a pre-trained VQ-Diffusion model  [^1] https://huggingface.co/microsoft/vq-diffusion-ithq that is trained on the ITHQ dataset (Tang et al., 2022). In all experiments, we optimize the parameters $\bm{\alpha}_{t}$ of the variational categorical distribution within the G2D2 algorithm’s optimization step using the Adam optimizer Kingma (2014). To balance the prior and likelihood terms in the objective function, we introduce hyperparameters. For the image inverse problem experiments, we used text prompts for the VQ-Diffusion model: “a photo of [Class Name]” for ImageNet and “a high-quality headshot of a person” for FFHQ.
@@ 4.2 Image inverse problem solving on ImageNet and FFHQ
Figure 5 shows the qualitative results of image inverse problem solving, and Tables 1 and 2 list the quantitative results. G2D2 performs comparably to the other methods using diffusion models trained in the continuous domain. Note that the pre-trained models used for each method are different, which particularly contributes to the superiority of pixel-domain methods on FFHQ. With DDRM, it is assumed that the amount of measurement noise is known and require the singular value decomposition of the linear operator. We also show images in the intermediate phase of the G2D2 algorithm in Figure 4.
Table 1: Quantitative evaluation on ImageNet 256$\times$256. Performance comparison of different methods on various linear tasks in image domain. Values show the mean over 100 images.
@@
Pixel-domain DPS (Chung et al., 2023) 0.367 22.61 0.443 19.04
DDRM (Kawar et al., 2022) 0.352 24.00 0.246 27.30
LDM PSLD (Rout et al., 2023) 0.332 24.43 0.365 24.04
ReSample (Song et al., 2024) 0.382 22.63 0.438 22.32
Discrete G2D2 (proposed) 0.349 23.20 0.375 22.71
G2D2 w/ Markov noise process 0.409 21.48 0.431 21.78
Figure 5: Qualitative results of G2D2 and DPS.
Table 2: Quantitative evaluation on FFHQ 256$\times$256. Performance comparison of different methods on various linear tasks in image domain. Values show mean over 100 images.
Prior Type Method SR ($\times$4) Gaussian deblurring
@@
Pixel-domain DPS (Chung et al., 2023) 0.227 26.73 0.225 26.02
DDRM (Kawar et al., 2022) 0.242 28.23 0.201 31.12
LDM PSLD (Rout et al., 2023) 0.276 27.62 0.304 27.37
ReSample (Song et al., 2024) 0.507 22.98 0.329 25.69
Discrete G2D2 (proposed) 0.271 26.93 0.287 26.35
G2D2 w/ Markov noise process 0.395 23.94 0.365 25.16
4.3 Ablation study on Graphical Models
It is possible to derive a similar algorithm to G2D2 that uses a Markov noise process as the graphical model. However, as discussed at the beginning of Section 3, this graphical model does not allow for the “re-masking” operation, which means it cannot correct errors that occur early in the sampling process. We refer to this variant as G2D2 w/ Markov noise process, and its performance is presented in Tables 1 and 2 on ImageNet and FFHQ, respectively. Additional qualitative results are provided in the Appendix C.3. The results indicate that the introduction of the star-shaped noise process significantly improves performance, making G2D2 comparable to continuous-based methods.
Figure 6: Results of executing G2D2 using a motion generation model as a prior in the motion inverse problem (path following generation task).
4.4 Motion inverse problem solving
As discussed in Section 3.4, our method can also be applied to Masked generative models. We conduct experiments to manipulate Generative Masked Motion Model (MMM) Pinyoanuntapong et al. (2024), a generative model for motion data, using gradient guidance. Specifically, we perform a path following task where generation is conditioned on the position information of the hip joint. Since joint position information can be calculated from motion data, this can also be treated within the framework of inverse problems. While there have been examples of achieving path following in motion generation models with continuous latent spaces (Song et al., 2023b; Uchida et al., 2024), we are the first to accomplish this using a motion generation model with discrete latent variables in a training-free manner. Appendix C.9 provides additional samples and detailed experimental information.
5 Conclusion
We proposed G2D2 for solving inverse problems using discrete diffusion models as priors. We demonstrated that G2D2 effectively addresses the limitation of discrete diffusion in inverse problem-solving by using a continuous relaxation technique and star-shaped noise process. Specifically, G2D2 approximates the posterior in inverse problems by optimizing the parameters of a variational distribution, composed of parameterized categorical distributions, at each time step of the diffusion process. Our experiments show that G2D2 performs comparable to its continuous counterparts, opening up possibilities for training-free applications of discrete diffusion models across a wide range of tasks.
Limitations and future works
G2D2 does not significantly surpass its continuous counterparts in terms of computational speed or performance. We anticipate that these limitations can be mitigated through the optimization of efficiency and the enhancement of prior models. The application to more complex problem settings, including nonlinear inverse problems, as well as to other domains such as audio and video, constitutes future work.
Ethics statement
Reproducibility statement
We will provide as detailed a description as possible regarding the reproduction of experiments in the Appendix, and we plan to release our code when this paper is published.
Appendix A Related work
In this section, we review the relevant prior works.
A.1 Leveraging Diffusion Models as Prior Models for Inverse Problems
Pixel-Domain Diffusion Models for Inverse Problems
Several methods have been proposed that utilize pixel-domain diffusion models for solving inverse problems. DDRM and DDNM (Kawar et al., 2022; Wang et al., 2023) assume linear operators and known noise levels, leveraging the singular value decomposition (SVD) of these operators. $\Pi$GDM (Song et al., 2023a) can handle certain classes of non-linear operators, such as low dynamic range, where a pseudo-inverse operator can be defined. Notably, $\Pi$GDM does not require SVD or gradient computations for such a case.
DPS (Chung et al., 2023) broadens the applicability to cases where operator gradients can be computed, enabling it to handle both linear and non-linear operators like phase retrieval and non-linear blur. Other notable methods in this category include RePaint (Lugmayr et al., 2022) and RED-Diff (Mardani et al., 2024).
@@ A.1 Leveraging Diffusion Models as Prior Models for Inverse Problems
A.2 Conditional Generation Using Discrete Diffusion Models as Priors

### v2

@@ 3.3 Variational approximation for feasible inference algorithm
7:    Initialize: $\bm{\alpha}_{t}\propto\exp(\gamma\log\bm{\alpha}_{t+1}+(1-\gamma)\log\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t}))$
@@
10:  $\bm{\alpha}_{t}=\operatorname*{arg\,min}_{\bm{\alpha}_{t}}D_{\mathrm{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})\right)-\mathbb{E}_{\mathbf{z}_{0}\sim\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}\left[\log q_{\text{star}}(\mathbf{y}|\mathbf{z}_{0})\right]$
11:  Sample $\mathbf{z}_{t-1}\sim p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})=\sum_{\mathbf{z}_{0}}q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{0})\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$
@@ 3.4 Implementation considerations
We use the Gumbel-Softmax trick (Jang et al., 2016; Maddison et al., 2016) to make the computation of the second term in (8) differentiable. At time step $t$, this process begins by generating Gumbel-Softmax samples using parameters of $\tilde{p}_{\bm{\alpha}}$ as follows: $\hat{z}_{0,i,k}=\mathrm{softmax}\left(\left(\log\alpha_{t,i,k}+g_{i,k}\right)/\tau\right)$, where $g_{i,k}$ are i.i.d. samples drawn from the Gumbel distribution, and $\tau$ (>0>0) is the temperature parameter. This procedure generates a “soft” categorical sample for each dimension in $\mathbf{z}_{0}$, indicating the proportional selection of each codebook element. As these proportions correspond to the contribution rate of each codebook element, we construct $\mathbf{Z}_{\text{Gumbel}}\in\mathbb{R}^{d_{\mathbf{z}}\times d_{b}}$ as their weighted sum: $(\mathbf{Z}_{\text{Gumbel}})_{i}=\sum_{k=1}^{K}\hat{z}_{0,i,k}\mathbf{b}_{k}$. Finally, we pass $\mathbf{Z}_{\text{Gumbel}}$ through the decoder to obtain the image $\mathbf{x}_{0}=D(\mathbf{Z}_{\text{Gumbel}})$. By substituting this image into the likelihood function $q_{\text{star}}(\mathbf{y}|\mathbf{x}_{0})$, we obtain the differentiable objective with respect to the variational parameter $\bm{\alpha}_{t}$, enabling continuous optimization. For linear inverse problems, the objective function will include the term $\|\mathbf{y}-\mathbf{A}\mathbf{x}_{0}(\bm{\alpha}_{t})\|_{2}^{2}$, excluding the constant term derived from measurement noise.
@@
At time step $t$, we are required to optimize the variational parameter $\bm{\alpha}_{t}$. To expedite this process, we can leverage the optimized values from the previous time step as the initialization for the optimization process, effectively reducing the number of required optimization steps. To achieve this, we introduce a forgetting coefficient $\gamma$ (where $0\leq\gamma\leq 1$) and initialize $\bm{\alpha}_{t}$ through a weighted sum of the previous optimized variables and the prior model’s output in the logarithm domain, given by $\bm{\alpha}_{t}\propto\exp(\gamma\log\bm{\alpha}_{t+1}+(1-\gamma)\log\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t}))$. The effectiveness of this strategy is discussed in Appendix E.6.
@@ 3.5 Application of G2D2 to masked generative models
As discussed in (Zheng et al., 2024), mask-absorbing discrete diffusion models and masked generative models, such as MaskGIT (Chang et al., 2022), share a similar framework. Apart from temporal conditioning, these models are nearly identical and are trained to approximate $q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t})$. Therefore, G2D2 can be straightforwardly applied to masked generative models. We empirically show this by providing an example of solving inverse problems using a masked generative model as a prior for motion data in the experimental section.
@@ 4.1 Experimental setup
We evaluate G2D2 on inverse problems in the image processing and compare it with other diffusion model-based inverse-problem-solving methods. We also demonstrate gradient-based guidance on a discrete-latent variable-based motion-domain generative model without additional training, showing the applicability of G2D2 to other domains.
@@
We conduct experiments on two tasks: (1) super-resolution (SR) and (2) Gaussian deblurring. For the SR task, the linear forward operator downscales the image by a factor of 4 using a bicubic resizer. For the Gaussian deblurring task, we set the kernel size to $61\times 61$ with a Gaussian kernel standard deviation of 3.0. The measurements are obtained by applying the forward operator to the ground truth images normalized to the range [1,1][-1,1], followed by the addition of Gaussian noise with a standard deviation of $0.05$. As metrics, we use the learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018) score to measure perceptual proximity to the original image, and the peak signal-to-noise ratio (PSNR) to measure the closeness of the signal.
@@
Following previous studies, we use the ImageNet (Deng et al., 2009) and Flickr-Faces-HQ (FFHQ) (Karras et al., 2019) datasets. The images are $\times256$. For comparison, we use a subset of 1000 images from each validation set.
@@
We compare DPS (Chung et al., 2023b), DDRM (Kawar et al., 2022), which use diffusion models trained in the pixel domain, and PSLD (Rout et al., 2023) and ReSample (Song et al., 2024), which use diffusion models trained in the latent space acquired from VAE (latent diffusion models) as baselines with G2D2.
@@
Regarding G2D2, for both the ImageNet and FFHQ experiments, we use a pre-trained VQ-Diffusion model  [^1] https://huggingface.co/microsoft/vq-diffusion-ithq that is trained on the ITHQ dataset (Tang et al., 2022). In all experiments, we optimize the parameters $\bm{\alpha}_{t}$ of the variational categorical distribution within the G2D2 algorithm’s optimization step using the RAdam optimizer (Liu et al., 2020). To balance the prior and likelihood terms in the objective function, we introduce hyperparameters. For the image inverse problem experiments, we used text prompts for the VQ-Diffusion model: “a photo of [Class Name]” for ImageNet and “a high-quality headshot of a person” for FFHQ. All experiments are performed on one RTX 3090 (24 GiB), but G2D2 itself never exceeded 4.7 GiB of VRAM and required 194 s per ImageNet image (Table 5 in the Appendix). Hence the full pipeline can be executed on widely available 8 GiB cards.
@@ 4.2 Image inverse problem solving on ImageNet and FFHQ
Figure 5 shows the qualitative results of image inverse problem solving, and Tables 1 and 2 list the quantitative results. As shown in Tables 1 and 2, G2D2 performs comparably to the methods using diffusion models trained in continuous domains. Notably, G2D2 achieves this performance while consuming significantly lower GPU memory (4.7GiB compared to 10.7GiB for DPS and 20.9GiB for PSLD, as detailed in Table 5 in the Appendix) and maintaining competitive computational speed among gradient-based methods. Note that the pre-trained models used for each method are different, which particularly contributes to the superiority of pixel-domain methods on FFHQ. With DDRM, it is assumed that the amount of measurement noise is known, and it requires the singular value decomposition of the linear operator. We also show images in the intermediate steps of the G2D2 algorithm in Figure 4.
Table 1: Quantitative evaluation on ImageNet $\times256$. Performance comparison of different methods on various linear tasks in image domain. Values show the mean over 1000 images.
@@
LPIPS($\downarrow$) PSNR($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$)
Pixel-domain DPS (Chung et al., 2023b) 0.362 22.67 0.432 19.49
DDRM (Kawar et al., 2022) 0.351 24.27 0.250 27.61
LDM PSLD (Rout et al., 2023) 0.331 24.02 0.359 23.95
ReSample (Song et al., 2024) 0.373 23.31 0.425 23.07
Discrete G2D2 (proposed) 0.340 23.82 0.367 23.37
G2D2 w/ Markov noise process 0.442 22.01 0.424 22.36
Figure 5: Qualitative results of G2D2 and DPS.
Table 2: Quantitative evaluation on FFHQ $\times256$. Performance comparison of different methods on various linear tasks in image domain. Values show the mean over 1000 images.
Prior Type Method SR ($\times$4) Gaussian Deblurring
LPIPS($\downarrow$) PSNR($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$)
Pixel-domain DPS (Chung et al., 2023b) 0.238 26.07 0.234 25.47
DDRM (Kawar et al., 2022) 0.252 28.09 0.209 30.89
LDM PSLD (Rout et al., 2023) 0.282 27.12 0.307 26.96
ReSample (Song et al., 2024) 0.508 23.07 0.336 25.91
Discrete G2D2 (proposed) 0.265 27.29 0.280 26.91
G2D2 w/ Markov noise process 0.369 25.15 0.340 25.96
4.3 Ablation study on Graphical Models
It is possible to derive a similar algorithm to G2D2 that uses a Markov noise process as the graphical model. However, as discussed at the beginning of Section 3, this graphical model does not allow for the “re-masking” operation, which means it cannot correct errors that occur early in the sampling process. We refer to this variant as G2D2 w/ Markov noise process, and its performance is presented in Tables 1 and 2 on ImageNet and FFHQ, respectively. Additional qualitative results are provided in the Appendix E.3. The results indicate that the introduction of the star-shaped noise process significantly improves performance, making G2D2 comparable to methods based on continuous diffusion.
4.4 Motion inverse problem solving
As discussed in Section 3.5, our method can also be applied to masked generative models. We conduct experiments to manipulate the Generative Masked Motion Model (MMM) (Pinyoanuntapong et al., 2024) using gradient guidance for a path-following task where generation is conditioned on hip joint position information. Since joint positions can be calculated from motion data, this fits within the inverse problems framework. While path following has been achieved with continuous latent space models (Song et al., 2023b; Uchida et al., 2024), we are the first to accomplish this using a discrete latent variable motion model without additional training. The Appendix E.13 provides additional samples and experimental details.
We compare G2D2 with OmniControl (Xie et al., 2024) and Guided Motion Diffusion (GMD) (Karunratanakul et al., 2023) on controllable motion generation. Following OmniControl’s setup, we evaluate using the HumanML3D test set with sparse conditioning (5 frames out of 196) on metrics including: FID, R-Precision, Diversity, Foot Skating ratio, Trajectory error (50cm), Location error (50cm), and Average error.
Table 3 presents the results. While OmniControl and GMD are specifically fine-tuned for this task, G2D2 requires no additional training yet achieves good FID and foot skating scores. However, OmniControl and GMD demonstrate better trajectory and location errors, indicating room for improvement. Note that G2D2 could potentially be combined with fine-tuned approaches (as OmniControl does) and enhanced with techniques like FreeDoM’s time-traveling method (Yu et al., 2023) to further improve performance.
Method FID R-prec. Diversity Foot Traj. Err Loc. err. Avg.
@@ 4.4 Motion inverse problem solving
G2D2 0.248 0.770 9.381 0.048 0.272 0.116 0.230
OmniControl (Xie et al., 2024) 0.278 0.705 9.582 0.058 0.053 0.015 0.043
GMD (Karunratanakul et al., 2023) 0.523 0.599 N/A 0.086 0.176 0.049 0.139
Table 3: Comparison of methods for controllable motion generation.
5 Conclusion
We proposed G2D2 for solving inverse problems using discrete diffusion models as priors. We demonstrated that G2D2 effectively addresses the limitations of discrete diffusion in inverse problem-solving by using a continuous relaxation technique and star-shaped noise process. Specifically, G2D2 approximates the posterior in inverse problems by optimizing the parameters of a variational distribution, composed of parameterized categorical distributions, at each time step of the diffusion process. Our experiments show that G2D2 performs comparably to its continuous counterparts, opening up possibilities for training-free applications of discrete diffusion models across a wide range of tasks.
Limitations and future work
While G2D2 already matches the image quality of continuous counterparts, its key advantage is an order‑of‑magnitude reduction in GPU memory footprint, enabling inference on consumer‑grade 8 GiB cards. Future work will focus on further accelerating sampling speed. We expect these gaps will narrow through efficiency optimizations and stronger prior models. Future work will also explore more challenging settings, such as nonlinear inverse problems, and extend G2D2 to additional modalities, including audio and video.
Acknowledgments
We are grateful to our colleague Satoshi Hayakawa for providing valuable feedback prior to submission, and we also thank the area chair and the reviewers for their constructive comments.
Appendix A Ethics statement
Appendix B Reproducibility statement
We provide a detailed description of the experimental reproduction in the Appendix, and our code is publicly available at https://github.com/sony/g2d2.
Appendix C Related work
In this section, we review the relevant prior works.
C.1 Leveraging Diffusion Models as Prior Models for Inverse Problems
Pixel-Domain Diffusion Models for Inverse Problems
@@ C.1 Leveraging Diffusion Models as Prior Models for Inverse Problems
DPS (Chung et al., 2023b) broadens the applicability to cases where operator gradients can be computed, enabling it to handle both linear and non-linear operators like phase retrieval and non-linear blur. Other notable methods in this category include RePaint (Lugmayr et al., 2022) and RED-Diff (Mardani et al., 2024).
@@
Application of Inverse Problem Solving in Various Domains
Solving inverse problems using diffusion models has enabled various real-world applications. In the image domain, diffusion models have been extensively studied and applied to tasks such as image deblurring, super-resolution, and inpainting (Lugmayr et al., 2022; Chung et al., 2023a; Zhu et al., 2023). In the audio domain, methods such as those proposed by Song et al. (2021a), Chung et al. (2023c), and Bian et al. (2024) have been developed to address tasks like dereverberation and audio restoration. Similarly, in the medical imaging domain, approaches like those introduced by Song et al. (2021a), Chung et al. (2023c), and Bian et al. (2024) have been used to improve image reconstruction and enhance diagnostic accuracy. These advancements demonstrate the versatility and effectiveness of diffusion models across different domains.
C.2 Conditional Generation Using Discrete Diffusion Models as Priors